# Towards Vision-Language Geo-Foundation Model: A Survey

### v1

@@ Abstract
Vision-Language Foundation Models (VLFMs) have made remarkable progress on various multimodal tasks, such as image captioning, image-text retrieval, visual question answering, and visual grounding. However, most methods rely on training with general image datasets, and the lack of geospatial data leads to poor performance on earth observation. Numerous geospatial image-text pair datasets and VLFMs fine-tuned on them have been proposed recently. These new approaches aim to leverage large-scale, multimodal geospatial data to build versatile intelligent models with diverse geo-perceptive capabilities, which we refer to as Vision-Language Geo-Foundation Models (VLGFMs). This paper thoroughly reviews VLGFMs, summarizing and analyzing recent developments in the field. In particular, we introduce the background and motivation behind the rise of VLGFMs, highlighting their unique research significance. Then, we systematically summarize the core technologies employed in VLGFMs, including data construction, model architectures, and applications of various multimodal geospatial tasks. Finally, we conclude with insights, issues, and discussions regarding future research directions. To the best of our knowledge, this is the first comprehensive literature review of VLGFMs. We keep tracing related works at https://github.com/zytx121/Awesome-VLGFM.
Index Terms:
Vision-language model, remote sensing foundation model, multimodal large language model
@@ 1 Introduction
Over the past decade, researchers have witnessed significant advancements in nearly all geospatial tasks, such as scene classification [1], object detection [2, 3], change detection [4], denoising [5], land use segmentation [6], disaster management [7], and geospatial localization [8], driven by deep learning and other artificial intelligence technologies. However, these models are specifically designed and trained for particular tasks, making them difficult to apply directly to other tasks. Even for the similar task, these models often exhibit poor generalization capabilities.
@@
For example, remote sensing object detection is one of the core tasks for earth observation. It requires manually annotating the position and category of each object, which is a time-consuming and labour-intensive process. Remote sensing images (RSIs) are captured from an overhead perspective by spaceborne or airborne sensors, which present unique viewpoints compared to natural images, leading to the development of oriented object detection. Since this task uses rotated bounding boxes to represent objects, remote sensing datasets annotated with rotated bounding boxes, such as DOTA [9], are required to support its training. Furthermore, the model architecture [10], loss functions [11], post-processing functions, and acceleration operators [12] must be modified based on standard object detection [13]. From this perspective, the application scenarios of specific models seem quite limited, lacking the ability to generalize across different tasks, even between two similar tasks.
@@
To reduce the resource wastage associated with training specific models from scratch for each task, foundational models [14] have emerged. These models are pre-trained on large-scale imagery, allowing them to handle various vision tasks with only fine-tuning small-scale custom datasets. In the field of remote sensing, prior studies on purely visual foundational models reveal immense potential for a generic model for earth observation, which is called Vision Geo-Foundation Models (VGFMs). VGFMs demonstrate remarkable generalization capabilities on a thorough evaluation, from single- to multi-modal, and static to temporal [15]. Although these models exhibit strong perceptual capabilities, they lack the ability to perform reasoning like humans. For instance, without corresponding annotated samples to support VGFM training, it cannot determine the specific function of a building in a remote sensing image by considering the surrounding environment and common sense, whereas humans can. Similarly, without annotated samples, VGFM cannot identify the brand or model of the car based on its features in aerial imagery, whereas humans can.
@@
Recently, the evolution of Large Language Models (LLMs) has revolutionized human-machine interaction. LLMs like BERT [16] exploit vast quantities of textual data to develop reasoning skills, showing substantial generalization capabilities across diverse tasks in natural language processing. However, LLMs process only discrete textual data and cannot handle images, while visual foundation models, though capable of processing image data, lack inferential abilities. The concept of vision-language foundation models (VLFMs) has been introduced to bridge these gaps. These innovative models are designed to perform perception and reasoning, integrating inputs from text and images. Since the release of GPT-4 Vision, the research on VLFMs has been fervently pursued, inspired by its impressive capabilities. Substantial efforts have been made in VLFMs, which are typically divided into contrastive, conversational, and generative paradigms. Below, we will briefly introduce the most influential works in three directions. Specifically, CLIP [17] adopts the contrastive paradigm, projecting visual and textual information into a unified representational space, thereby facilitating a bridge for downstream vision-language tasks. LLaVA [18] embodies the conversational paradigm, endowing LLMs with the capability for context-aware dialogue across textual and visual modalities. Stable Diffusion [19], epitomizing the generative paradigm, leverages deep learning to produce high-quality, detailed images from textual descriptions, thereby advancing the capabilities in image synthesis and creative visual applications.
@@
When VLFMs are applied to earth observation, this paper refers to them as Vision-Language Geo-Foundation Models (VLGFMs). To date, VLGFMs can also be categorized into contrastive, conversational, and generative types. Figure 1 lists the representative VLGFM developed and their release dates. It can be seen that VLGFMs first appeared in the second quarter of 2023. And now, the number of related works is in a period of rapid growth. Notably, current innovations in VLGFM are primarily focused on collecting training data, with relatively few modifications to the model architecture. Most efforts involve fine-tuning based on the LLaVA [18] and MiniGPT-4 [20] frameworks using custom-built remote sensing instruction-following sets.
@@
TABLE I: Model architectural of VLGFMs. Sort models by date of initial release. Abbreviations IS, IC, VC, IR, OC, REC, REG, RES, and GM in the table stand for image scene, image captioning, video captioning, image retrieval, object counting, referring expression comprehension, referring expression generation, referring expression segmentation, and geometric measurement, respectively. For each model, we indicate the LLM used in its best configuration as shown in the original paper ($\lozenge$: training from scratch; $\blacklozenge$: fine-tuning; $\blacktriangle$: fine-tuning with parameter-efficient fine-tuning techniques [21]; $\bigstar$: frozen).
@@
[29] ConvNeXt-L (CLIP) REC, REG
@@
[34] OVD, REC, REG
TABLE II: Evaluate data of VLGFMs. Abbreviations IC, VQA, VG, IS, and IR in the table stand for image captioning, visual question answering, visual grounding, image scene classification, and image retrieval, respectively (If more than 1 models use the dataset to evaluate a task, it is highlighted in blue. If more than 2 models use it, it is highlighted in red).
@@
[24] Sydney-Captions RSVQA-HR
@@
[27] AID, fMoW, RESISC45, RSICD,
@@
[28] Sydney-Captions, RSVQA-HR DIOR-RSVG
@@
[29] RSVQA-HR MAR20 CLRS, NaSC-TG2
@@
[30] RSVQA-HR DIOR-RSVG SIRI-WHU, EuroSAT,
@@
[31] Sydney-Captions, UCM
@@
H2RSVLM - RSVQA-LR DIOR-RSVG NWPU, METER-ML, - RSSA
[32] RSVQA-HR SIRI-WHU, AID, HqDC-Instruct
@@
[34] RSVQA-HR SIRI-WHU FIT-RSRC
@@ 2.1 Concept Definition
Foundation models is a kind of model with broad applicability and general capabilities after training on large-scale data. These models typically have a large number of parameters and strong learning abilities, enabling them to excel in a variety of tasks. Although foundation models may not be ubiquitous at present, they appear poised to become the basis of widespread technological innovations and exhibit the key hallmarks of a general-purpose technology [35]. The field surveyed in this paper can handle multiple tasks simultaneously, and vision-language models that support only a single task are outside the scope of our research. Please refer to [14] for those interested in task-specific vision-language models.
Geo-foundation models are a class of models specifically designed for processing geospatial data through visual information. These models utilize various types of geospatial visual data, such as remote sensing imagery, satellite photos, and aerial images, to perform detailed analysis and support a range of geographic applications. Although this research is currently a popular topic, it is not the focus of this paper. For those interested in purely visual geo-foundation models, please refer to [36] for further information.
@@ 2.2 History and Roadmap
Before introducing the VLGFM details, it is necessary to review its progress. In Figure 1, we summarize the timeline of VLGFM and list some representative works. The development of VLFMs in the remote sensing domain started relatively late. It was not until the emergence of the works LLaVA [18] and MiniGPT-4 [20] in 2023 that VLGFM-related work began to take shape. Most existing conversational VLGFMs have been implemented based on the open-source frameworks released by these two works. Since June 2023, the first batch of VLGFMs began to emerge. Among them, RemoteCLIP [22] was the first contrastive VLGFM, supporting image scene classification and image-text retrieval tasks. RSGPT [24] was the first conversational VLGFM, supporting image captioning and visual question answering tasks. RS5M [23] was the first million-scale remote sensing image-text pair dataset made publicly available. Since October 2023, the number of VLGFM works has steadily increased. Among them, DiffusionSat [37] is the first high-quality text-to-image generation work in the remote sensing field. It can control the generation of RSIs using conditions such as latitude and longitude, season, cloud cover, and ground sampling distance (GSD). GeoChat [25] is the first conversational VLGFM to support visual grounding task., while SkyEyeGPT [28] is the first to support remote sensing video captioning task. H2RSVLM [32] is the first to introduce the concept of trustworthiness, where the model refuses to answer uncertain questions. SkySenseGPT [34] introduces the concept of graphs to enhance VLGFM’s ability to perceive and reason relationships between remote sensing targets as well as its capability for complex comprehension. It is worth mentioning that research on VLGFMs is more akin to a data-centric project rather than a model-centric one. As seen from Table I, the structure of each model is quite similar, with no significant modifications. However, Tables II shows that the data selection for different VLGFMs varies considerably.
@@ 2.3 Tasks, Datasets, and Metrics
$\bullet$ Tasks. VLGFMs are capable of addressing a wide range of earth observation tasks, including image scene classification (IS) [38], image retrieval (IR) [39], image captioning (IC) [40], visual question answering (VQA) [41], and visual grounding (VG) [42]. Among these tasks, the VQA task is particularly special as it serves as a means to evaluate most multi-modal tasks in VLGFMs. Specifically, the inputs of VQA include images and text prompts, and the output is text. By using different text prompts, answers for various tasks can be obtained. VG is a more practical task in real-world applications and is mainly divided into two sub-tasks: referring expression comprehension and referring expression generation. Detailed introduction to the various capabilities of VLGFMs will be provided in Section 3.3.
$\bullet$ Datasets. In Table II, we list the datasets commonly used in experimental validation of existing VLGFMs and categorize them into three types based on usage frequency: recommended, common, and rare. Red, blue, and black fonts represent these categories. For the image captioning task, the recommended datasets are UCM-Captions and Sydney-Captions [40]; the commonly used datasets are RSICD [43] and NWPU-Captions [44]; and the rare datasets are CapERA [45] and UAV [46]. For the VQA task, the recommended datasets are RSVQA-LR and RSVQA-HR [41]. The recommended dataset for the visual grounding task is DIOR-RSVG [47]; the commonly used dataset is RSVG [42]; and the rare dataset is MAR20 [48]. For the image scene task, the recommended datasets are EuroSAT [49], RESISC45 [50], WHU-RS19 [51], SIRI-WHU[52], and AID [53]; the commonly used datasets are PatternNet [54], RSI-CB256 [55], METER-ML [56] and fMoW [57]. For the image retrieval task, the recommended datasets are RSITMD [39], RSICD [43], and UCM [58]. At the same time, some VLGFMs have also released their own evaluation sets constructed for different specific capabilities.
@@
Additionally, experts can be employed to score the outcomes of responses. For instance, RSGPT [24] utilizes experts to categorize the results of captioning tasks into four distinct levels. Since the advent of large language models, efforts have been made to leverage these models for scoring the outcomes of various vision-language multimodal tasks. For choice questions, the output of some models may not always meet the requirements due to imperfect instruction-following capabilities. Their responses may be completely unrelated to the questions, often failing to convey the intended meaning. Their responses may be completely unrelated to the questions, consisting of meaningless numbers and special symbols. MMBench [59] employs ChatGPT [60] as a judge to identify the option most similar to the model’s output and regard it as the model’s prediction result.

### v2

@@ Abstract
Vision-Language Foundation Models (VLFMs) have made remarkable progress on various multimodal tasks, such as image captioning, image-text retrieval, visual question answering, and visual grounding. However, most methods rely on training with general image datasets, and the lack of geospatial data leads to poor performance on earth observation. Numerous geospatial image-text pair datasets and VLFMs fine-tuned on them have been proposed recently. These new approaches aim to leverage large-scale, multimodal geospatial data to build versatile intelligent models with diverse geo-perceptive capabilities, which we refer to as Vision-Language Geo-Foundation Models (VLGFMs). This paper thoroughly reviews VLGFMs, summarizing and analyzing recent developments in the field. In particular, we introduce the background and motivation behind the rise of VLGFMs, highlighting their unique research significance. Then, we systematically summarize the core technologies employed in VLGFMs, including data construction, model architectures, and applications of various multimodal geospatial tasks. Finally, we conclude with insights, issues, and discussions regarding future research directions.
journal: Research Article
@@ 1 Introduction
Over the past decade, researchers have witnessed significant advancements in nearly all geospatial tasks, such as scene classification hongdanfeng, object detection 7560644; 9852282, change detection 8022932, denoising denoise, land use segmentation 8309343, disaster management 9460988, and geospatial localization 4587784, driven by deep learning and other artificial intelligence technologies. However, these models are specifically designed and trained for particular tasks, making them difficult to apply directly to other tasks. Even for the similar task, these models often exhibit poor generalization capabilities.
@@
For example, remote sensing object detection is one of the core tasks for earth observation. It requires manually annotating the position and category of each object, which is a time-consuming and labour-intensive process. Remote sensing images (RSIs) are captured from an overhead perspective by spaceborne or airborne sensors, which present unique viewpoints compared to natural images, leading to the development of oriented object detection. Since this task uses rotated bounding boxes to represent objects, remote sensing datasets annotated with rotated bounding boxes, such as DOTA Xia_2018_dota, are required to support its training. Furthermore, the model architecture Yang2021R3Det, loss functions Yang2021Rethinking, post-processing functions, and acceleration operators zhou2019iou must be modified based on standard object detection mmrotate. From this perspective, the application scenarios of specific models seem quite limited, lacking the ability to generalize across different tasks, even between two similar tasks.
@@
To reduce the resource waste associated with the training of specific models from scratch for each task, foundational models li2024visionlanguage have emerged. These models are pre-trained on large-scale imagery, allowing them to handle various vision tasks with only fine-tuning small-scale custom datasets. In the field of remote sensing, prior studies on purely visual foundational models reveal immense potential for a generic model for earth observation, which is called Vision Geo-Foundation Models (VGFMs). VGFMs demonstrate remarkable generalization capabilities in a thorough evaluation, from single- to multi-modal, and static to temporal guo2024skysense. Although these models exhibit strong perceptual capabilities, they lack the ability to perform reasoning like humans. For instance, without corresponding annotated samples to support VGFM training, it cannot determine the specific function of a building in a remote sensing image by considering the surrounding environment and common sense, whereas humans can. Similarly, without annotated samples, VGFM cannot identify the brand or model of the car based on its features in aerial imagery, whereas humans can.
@@
Recently, the evolution of Large Language Models (LLMs) has revolutionized human-machine interaction. LLMs like BERT devlin2019bert take advantage of large amounts of textual data to develop reasoning skills, showing substantial generalization capabilities on various tasks in natural language processing. However, LLMs process only discrete textual data and cannot handle images, while visual foundation models, though capable of processing image data, lack inferential abilities. The concept of vision-language foundation models (VLFMs) has been introduced to bridge these gaps. These innovative models are designed to perform perception and reasoning, integrating inputs from text and images. Since the release of GPT-4 Vision, the research on VLFMs has been fervently pursued, inspired by its impressive capabilities. Substantial efforts have been made in VLFMs, which are typically divided into contrastive, conversational, and generative paradigms. Below, we will briefly introduce the most influential works in three directions. Specifically, CLIP pmlr-v139-radford21a adopts the contrastive paradigm, projecting visual and textual information into a unified representational space, thereby facilitating a bridge for downstream vision-language tasks. LLaVA llava embodies the conversational paradigm, endowing LLMs with the capability to recognize context in textual and visual modalities. Stable Diffusion rombach2022highresolution, epitomizing the generative paradigm, leverages deep learning to produce high-quality, detailed images from textual descriptions, thereby advancing the capabilities in image synthesis and creative visual applications.
@@
When VLFMs are applied to earth observation, this paper refers to them as Vision-Language Geo-Foundation Models (VLGFMs). To date, VLGFMs can also be categorized into contrastive, conversational, and generative types. Figure 1 lists the representative VLGFM developed and their release dates. It can be seen that VLGFMs first appeared in the second quarter of 2023. And now, the number of related works is in a period of rapid growth. Notably, current innovations in VLGFM are primarily focused on collecting training data, with relatively few modifications to the model architecture. Most efforts involve fine-tuning based on the LLaVA llava and MiniGPT-4 zhu2023minigpt4 frameworks using custom-built remote sensing instruction-following sets.
@@
Table 1: Model architectural of VLGFMs. Sort models by date of initial release. Abbreviations IS, IC, VC, IR, OC, OVD, REC, REG, RES, and GM in the table stand for image scene, image captioning, video captioning, image retrieval, object counting, open-vocabulary detection, referring expression comprehension, referring expression generation, referring expression segmentation, and geometric measurement, respectively. For each model, we indicate the LLM used in its best configuration as shown in the original paper ($\lozenge$: training from scratch; $\blacklozenge$: fine-tuning; $\blacktriangle$: fine-tuning with parameter-efficient fine-tuning techniques si2024flora; $\bigstar$: frozen).
@@
zhang2024earthgpt ConvNeXt-L (CLIP) REC, REG
@@
luo2024skysensegpt OVD, REC, REG
RSUniVLM Qwen2-0.5B SigLIP-400M a two-layer MLP VQA, RES, 4*A40 (40G)
liu2024rsunivlm IS, IC, REC
EarthDial Phi-3-mini InternViT-300M a two-layer MLP VQA, OVD, 8*A100 (80G)
soni2024earthdial IS, IC, REC
RingMo-Agent DeepSeekMoE-3B SigLIP-SO400M-384 a two-layer MLP VQA, OVD, 8*A100 (80G)
hu2025ringmoagent IS, IC
RSCoVLM Qwen2.5-7B Modified ViT a two-layer MLP VQA, OVD, 8*A800 (80G)
li2025rscovlm IS, REC
GeoZero Qwen3-8B SigLIP2-SO-400M a two-layer MLP VQA, IC, 8*A100 (40G)
wang2025geozero IS, REC, VQA
Table 2: Evaluate data of VLGFMs. Abbreviations IC, VQA, VG, IS, and IR in the table stand for image captioning, visual question answering, visual grounding, image scene classification, and image retrieval, respectively (If more than 1 models use the dataset to evaluate a task, it is highlighted in blue. If more than 2 models use it, it is highlighted in red).
@@
hu2023rsgpt Sydney-Captions RSVQA-HR
@@
wang2024skyscript AID, fMoW, RESISC45, RSICD,
@@
zhan2024skyeyegpt Sydney-Captions, RSVQA-HR DIOR-RSVG
@@
zhang2024earthgpt RSVQA-HR MAR20 CLRS, NaSC-TG2
@@
muhtar2024lhrsbot RSVQA-HR DIOR-RSVG SIRI-WHU, EuroSAT,
@@
silva2024large Sydney-Captions, UCM
@@
H2RSVLM - RSVQA-LR DIOR-RSVG RESISC45, METER-ML, - RSSA
pang2024h2rsvlm RSVQA-HR SIRI-WHU, AID, HqDC-Instruct
@@
luo2024skysensegpt RSVQA-HR SIRI-WHU FIT-RSRC
RSUniVLM LEVIR-MCI RSVQA-LR DIOR-RSVG RESISC45, - WHU-CD,
liu2024rsunivlm RSVQA-HR VRSBench SIRI-WHU, AID, Vaihingen,
WHU-RS19 UDD5, VDD
RSVQA-LR, RSVG, GeoChat, AID, WHU-RS19,
RSCoVLM RSVQA-HR, DIOR-RSVG METER-ML, DOTA-v1.0
li2025rscovlm LRS-VQA, VRSBench RESISC45,
VRSBench AVVG UCM
UCM-Captions RSVG, AID, WHU-RS19,
GeoZero Sydney-Captions RSVQA-HR, DIOR-RSVG SIRI-WHU, CHOICE
wang2025geozero NWPU-Captions XLRSBench VRSBench RESISC45,
RSICD EuroSAT,UCM
@@ 2.1 Concept Definition
Foundation models is a kind of model with broad applicability and general capabilities after training on large-scale data. These models typically have a large number of parameters and strong learning abilities, enabling them to excel in a variety of tasks. Although foundation models may not be ubiquitous at present, they appear poised to become the basis of widespread technological innovations and exhibit the key hallmarks of a general-purpose technology bommasani2022opportunities. The field surveyed in this paper can handle multiple tasks simultaneously, and vision-language models that support only a single task are outside the scope of our research. Please refer to li2024visionlanguage for those interested in task-specific vision-language models.
Geo-foundation models are a class of models specifically designed for processing geospatial data through visual information. These models utilize various types of geospatial visual data, such as remote sensing imagery, satellite photos, and aerial images, to perform detailed analysis and support a range of geographic applications. Although this research is currently a popular topic, it is not the focus of this paper. For those interested in purely visual geo-foundation models, please refer to 10282966 for further information.
@@ 2.2 History and Roadmap
Before introducing the VLGFM details, it is necessary to review its progress. In Figure 1, we summarize the timeline of VLGFM and list some representative works. The development of VLFMs in the remote sensing domain started relatively late. It was not until the emergence of the works LLaVA llava and MiniGPT-4 zhu2023minigpt4 in 2023 that VLGFM-related work began to take shape. Most existing conversational VLGFMs have been implemented based on the open-source frameworks released by these two works. Since June 2023, the first batch of VLGFMs began to emerge. Among them, RemoteCLIP liu2024remoteclip was the first contrastive VLGFM, supporting image scene classification and image-text retrieval tasks. RSGPT hu2023rsgpt was the first conversational VLGFM, supporting image captioning and visual question answering tasks. RS5M zhang2024rs5m was the first million-scale remote sensing image-text pair dataset made publicly available. Since October 2023, the number of VLGFM works has steadily increased. Among them, DiffusionSat khanna2023diffusionsat is the first high-quality text-to-image generation work in the remote sensing field. It can control the generation of RSIs using conditions such as latitude and longitude, season, cloud cover, and ground sampling distance (GSD). GeoChat kuckreja2023geochat is the first conversational VLGFM to support visual grounding task., while SkyEyeGPT zhan2024skyeyegpt is the first to support remote sensing video captioning task. H2RSVLM pang2024h2rsvlm is the first to introduce the concept of trustworthiness, where the model refuses to answer uncertain questions. SkySenseGPT luo2024skysensegpt introduces the concept of graphs to enhance VLGFM’s ability to perceive and reason relationships between remote sensing targets as well as its capability for complex comprehension. It is worth mentioning that research on VLGFMs is more akin to a data-centric project rather than a model-centric one. As seen from Table 1, the structure of each model is quite similar, with no significant modifications. However, Tables 2 shows that the data selection for different VLGFMs varies considerably.
@@ 2.3 Tasks, Datasets, and Metrics
$\bullet$ Tasks. VLGFMs are capable of addressing a wide range of earth observation tasks, including image scene classification (IS) roberts2023satin, image retrieval (IR) rsitmd, image captioning (IC) sydneycaption, visual question answering (VQA) rsvqa, and visual grounding (VG) rsvg. Among these tasks, the VQA task is particularly special as it serves as a means to evaluate most multi-modal tasks in VLGFMs. Specifically, the inputs of VQA include images and text prompts, and the output is text. By using different text prompts, answers for various tasks can be obtained. VG is a more practical task in real-world applications and is mainly divided into two sub-tasks: referring expression comprehension and referring expression generation. Detailed introduction to the various capabilities of VLGFMs will be provided in Section 3.3.
$\bullet$ Datasets. In Table 2, we list the datasets commonly used in experimental validation of existing VLGFMs and categorize them into three types based on usage frequency: recommended, common, and rare. Red, blue, and black fonts represent these categories. For the image captioning task, the recommended datasets are UCM-Captions and Sydney-Captions sydneycaption; the commonly used datasets are RSICD rsicd and NWPU-Captions nwpucaption; and the rare datasets are CapERA capera and UAV uavcaption. For the VQA task, the recommended datasets are RSVQA-LR and RSVQA-HR rsvqa. The recommended dataset for the visual grounding task is DIOR-RSVG dior-rsvg; the commonly used dataset is RSVG rsvg; and the rare dataset is MAR20 wenqi2024mar20. For the image scene task, the recommended datasets are EuroSAT eurosat, RESISC45 resisc45, WHU-RS19 whu-rs19, SIRI-WHUsiri-whu, and AID aid-data; the commonly used datasets are PatternNet patternnet, RSI-CB256 rsi-cb, METER-ML zhu2022meterml and fMoW fmow. For the image retrieval task, the recommended datasets are RSITMD rsitmd, RSICD rsicd, and UCM ucm. At the same time, some VLGFMs have also released their own evaluation sets constructed for different specific capabilities.
@@
Additionally, experts can be employed to score the outcomes of responses. For instance, RSGPT hu2023rsgpt utilizes experts to categorize the results of captioning tasks into four distinct levels. Since the advent of large language models, efforts have been made to leverage these models for scoring the outcomes of various vision-language multimodal tasks. For choice questions, the output of some models may not always meet the requirements due to imperfect instruction-following capabilities. Their responses may be completely unrelated to the questions, often failing to convey the intended meaning. Their responses may be completely unrelated to the questions, consisting of meaningless numbers and special symbols. MMBench liu2023mmbench employs ChatGPT gpt3 as a judge to identify the option most similar to the model’s output and regard it as the model’s prediction result.